# Effect of the Large Magellanic Cloud on the kinematics of Milky Way satellites and virial mass estimate

### v1

@@ Abstract
We present a study illustrating the effects of the passage of a Large Magellanic Cloud (LMC) mass satellite on the distance and velocity distributions of satellites in $\Lambda+$Cold Dark Matter simulations of Milky Way (MW) sized halos. In agreement with previous studies, we find that during such a passage the velocity distribution develops a high-velocity tail due to the reflex motion of the inner part of the halo, which can bias velocity-based virial halo mass estimates. When the velocity distribution of MW satellites is corrected for effects of the LMC passage, it is consistent with the distributions in halos of masses as low as $M_{\rm 200c}=8\times 10^{11}\,\,M_{\odot}$ and as high as $1.5\times 10^{12}\,\,M_{\odot}$. We present a new halo mass estimator $M_{\rm 200c}=c\sigma^{2}_{\rm 3D}r_{\rm med}$, where $c$ is a constant calibrated using satellite systems in the simulated MW-sized halos, $\sigma^{2}_{\rm 3D}$ is the variance of 3D velocities taken with the sign of the radial velocity of each satellite, and $r_{\rm med}$ is the median halocentric distance of the satellites. We show that the estimator has only $s=8\%$ scatter around the median relation of the estimated and true halo masses and deviates by $<2s$ from the median during the pericentric passage of an LMC-like subhalo. This is because $\sigma^{2}_{\rm 3D}$ and $r_{\rm med}$ deviate in the opposite directions during such passages. We apply the estimator to the MW satellite system and estimate the virial mass of the Milky Way of $M_{\rm 200c}=9.96\pm 1.45\times 10^{11}\,\,M_{\odot}$, in good agreement with several recent estimates using other methods.
@@ 1 Introduction
The improved velocity accuracy and sophistication of the analyses led to reduced uncertainties in halo mass estimates and reduced discrepancies between different studies. Nevertheless, recent MW halo virial mass measurements still allow a fairly wide range of values $\approx 6-20\times 10^{11}\,M_{\odot}$ (e.g., Vasiliev et al., 2021; Correa Magnus & Vasiliev, 2022; Koposov et al., 2023; Zhou et al., 2023; Roche et al., 2024).
@@
Such a large LMC mass results in perturbations of orbits of individual MW satellites (Erkal et al., 2020; Pace et al., 2022; Makarov et al., 2023), a significant reflex motion of the inner Milky Way, and asymmetries of the potential in which dwarf satellites and orbit (Gómez et al., 2015; Erkal et al., 2020; Garavito-Camargo et al., 2019; Garavito-Camargo et al., 2021; D’Souza & Bell, 2022). These effects have been detected using different tracers in the Milky Way halo (Petersen & Peñarrubia, 2020, 2021; Erkal et al., 2021; Conroy et al., 2021). In particular, the reflex motion of the Milky Way affects the velocities of the outer satellites, such as Antlia II (Ji et al., 2021) and Leo I (Erkal et al., 2020). The use of the latter in particular tended to produce higher halo mass estimates in prior analyses. Erkal et al. (2020) also showed that the mass estimator of Watkins et al. (2010) applied to the velocities of satellites and stellar tracers overestimates the enclosed mass $M(<r)$ by up to $50\%$ at large $r$ if the LMC-induced reflex motion is neglected. The presence of the bound Magellanic Clouds pair also affects statistical estimates of the Milky Way halo mass (González et al., 2013). Finally, the approach of the LMC has a large effect on the spatial distribution of MW satellites (Nadler et al., 2020; Manwadkar & Kravtsov, 2022; Garavito-Camargo et al., 2023).
@@
Large effects of the LMC on the phase-space distribution of the potential Milky Way tracers need to be accounted for in theoretical inferences of halo mass that use these tracers. This is challenging, however, because only $\approx 5-10\%$ of the MW-sized halos have a pair of LMC-like satellite with a pericenter close to $\approx 50$ kpc (Boylan-Kolchin et al., 2010; Busha et al., 2011). Thus, large samples of halos need to be simulated with a high resolution sufficient to resolve and model satellite evolution reliably. Alternatively, the analyses can focus on the properties of tracers that are not significantly affected by the dynamical perturbations due to the LMC, such as specific angular momentum (Patel et al., 2018).
@@
In this study, we explore the effects of the LMC on the velocity distribution of satellite systems using a Caterpillar suite of high-resolution simulations of MW-sized halos (Griffen et al., 2016, see Section 2). We demonstrate that similarly to the real LMC, LMC-like objects in simulations induce a reflex motion onto their host halo that creates a high-velocity tail in the distribution of halocentric 3D satellite velocities (Section 3.1). We also show how accounting for this tail affects the conclusion about MW halo mass based on the velocity distribution. We then present a new halo mass estimator that has only $\approx 8\%$ scatter around the true mass when applied to simulated halos (Section 3.2) and which is affected by the LMC-like objects by only $\lesssim 10\%$. We use the estimator calibrated with the simulated halos to infer the MW halo mass of $M_{\rm 200c}=9.93\pm 1.45\times 10^{11}\,M_{\odot}$. We summarize our results and conclusions in Section 4.
@@ 2 Modeling Milky Way satellite system
The mass evolution tracks of subhaloes of MW-sized host haloes are used as input for the GRUMPY galaxy formation model, which evolves various properties of gas and stars of the galaxies they host. It is a regulator-type galaxy formation framework (e.g., Krumholz & Dekel, 2012; Lilly et al., 2013; Feldmann, 2013) designed to model galaxies of $\lesssim L_{\star}$ luminosity (Kravtsov & Manwadkar, 2022) which follows the evolution of several key galaxy properties by solving a system of coupled different equations. The model accounts for UV heating after reionization and associated gas accretion suppression onto small mass haloes, galactic outflows, a model for gaseous disk and its size, molecular hydrogen mass, star formation, etc. The evolution of the half-mass radius of the stellar distribution is also modeled. The galaxy model parameters used in this study are identical to those used in Manwadkar & Kravtsov (2022). Importantly for this study, the model was shown to reproduce the luminosity function and radial distribution of the Milky Way satellites and the size-luminosity relation of observed dwarf galaxies (Manwadkar & Kravtsov, 2022). Here we use the model to predict luminosities and stellar masses of satellite galaxies around the MW-sized haloes from the Caterpillar suite.
@@
Figure 3: The cumulative distribution of 3D velocities of satellites in the Milky Way and simulated halos of different halo mass ranges. The velocity distribution of the Milky Way satellites is shown by the thick solid line. The long-dashed, dot-dashed, and dashed lines show velocity CDFs of satellites in the low, intermediate, and high mass host halo mass ranges. The satellites have been selected using the distance matching procedure described in the text (Section 3.1).
Figure 4: Top panel: evolution of distance (in physical kpc) of two LMC-like satellites to the center of their host halo from $z=0.5$ to $z=0$ in two different Caterpillar host halos. The satellites are on their first infall and reach their pericenter of $\approx 40-50$ kpc at $z_{\rm LMC,peri}\approx 0.04-0.05$. Bottom panel: Evolution of halo masses (solid lines) and stellar masses (dashed lines) of the two LMC-like satellites shown in the top panel. The satellites reach peak halo masses of $M_{\rm peak}\approx 1.16\times 10^{11}\,M_{\odot}$ and $M_{\rm peak}\approx 1.35\times 10^{11}\,M_{\odot}$ at $z\approx 0.1$ before they start to lose mass due to tidal stripping. The stellar mass is computed using our galaxy formation model and continues to grow due to ongoing star formation and reaches $M_{\star}\approx 3\times 10^{9}\,M_{\odot}$, comparable to the stellar mass of the observed LMC.
@@
3 Results
3.1 Effect of the LMC on the velocity distribution of satellites
Comparison to the Milky Way requires a satellite population whose cumulative Galactocentric distance distribution matches that of the MW satellites. To obtain this, we draw a set of random distances from the MW satellite distance distribution using the inverse transform sampling method and spline approximation of the inverse cumulative distribution function of MW satellite distances. We then selected subhalos in each simulated MW-sized halo that were closest to each of the drawn distances. Figure 1 shows that this distance-matching technique produces a satellite population with cumulative distance distribution consistent with the distance distribution of MW the satellites. At the same time, Figure 2 demonstrates that the 3D velocity distribution of the distance-matched subhalos is statistically consistent with the velocity distribution of Milky Way satellites.
Hosts 4 and 53 from the Caterpillar host halo tracks include the passage of subhalos that reach peak masses of $M_{\rm peak}=1.16\times 10^{11}\,\,M_{\odot}$ and $M_{\rm peak}=1.35\times 10^{11}\,\,M_{\odot}$, respectively. These are comparable to the mass of the LMC with $M_{\rm peak}=1.3\pm 0.3\times 10^{11}\,M_{\odot}$ (Erkal et al., 2019; Vasiliev et al., 2021; Shipp et al., 2021; Watkins et al., 2024). The time evolution of the halocentric distance and mass of these subhalos are shown in Figure 4. The satellites reach their peak masses $M_{\rm peak}\approx 1.16\times 10^{11}\,M_{\odot}$ and $M_{\rm peak}\approx 1.35\times 10^{11}\,M_{\odot}$ at $z\approx 0.1$ and begin to lose mass due to tidal stripping during their pericentric passage at $z\approx 0.05$. As the LMC-like satellites continue to be tidally stripped, stellar mass steadily increases due to continued star formation and reaches a peak of $M_{\star}\approx 3\times 10^{9}\,M_{\odot}$, comparable to the observational estimates of the LMC’s stellar mass.
@@ 3.1 Effect of the LMC on the velocity distribution of satellites
For one of the host halos, Figure 5 shows the satellite velocity distribution at $z=0$ and at the time of pericentric passage $z\approx 0.05$, as well as the 15.9 - 84.1 and 2.3 - 97.7 percentile intervals of the velocity distribution in bins of distance. The effect of the LMC-like satellite is greatest near the pericenter, where the subhalo velocity distribution becomes skewed toward larger negative velocities. At $z=0$, the velocities return to a symmetric distribution. The effect of the LMC-like subhalo passage on the cumulative velocity distribution is shown in Figure 6, which shows that the satellite distribution has a tail of higher velocities near the pericentric passage. The cumulative velocity distribution at $z=0$, on the other hand, is similar to the distribution before the LMC-like subhalo entered the host halo at $z=0.1$. The lower panel of Figure 5 shows that the cumulative distance distribution of satellites becomes more centrally concentrated during the pericentric passage of the LMC-like subhalo. Like the velocity distribution, the distance distribution at $z=0$ becomes closer to the distribution before the passage at $z=0.1$.
@@
The physical origin of the large change in the velocity and distance distribution of satellites during the pericentric passage of an LMC-like subhalo is the reflex motion of the central region of the MW-sized host halo in response to the gravitational pull of the massive subhalo. During the passage, the mean motion of the host halo relative to the subhalo system is $\approx 40-50\,\rm km/s$, comparable to the measured magnitude of the Milky Way disk reflex motion (e.g., Petersen & Peñarrubia, 2021).
@@
The large effect of the LMC on the velocity and distance distribution needs to be taken into account in the estimates of MW halo mass (Erkal et al., 2020). Although the effect becomes small well after the passage, the LMC in the Milky Way has just passed its pericenter and therefore effects on the MW satellite velocities are expected to be significant. To illustrate the effect of change in velocity distribution on halo mass inference, we approximate the difference in the velocity distributions at the pericentric passage of the LMC-like subhalo and at $z=0$ as a function of 3D velocity and apply the difference as the correction to the velocity of each observed satellite of the Milky Way.
@@
Patel et al. (2018) used the specific angular momenta of satellites to estimate MW halo mass and showed that they are not sensitive to deviations from dynamic equilibrium. The upper panel of Figure 9 shows the cumulative distribution of specific angular momentum $j=|\mathbf{r}\times\mathbf{v}_{\rm 3D}|$ at three different epochs, $z=0.1$, $z=z_{\rm LMC,peri}$ and $z=0$, for one of the host halos with an LMC-like satellite passing through its pericenter at $z_{\rm LMC,peri}\approx 0.05$. The effect of the LMC-like satellite on the $j$ distribution is indeed very small, which is consistent with the conclusion of Patel et al. (2018). Thus, the use of the angular momentum to estimate halo mass would not require a correction for the effect of the LMC.
At the same time, the lower panel of Figure 9 shows that the cumulative $j$ distributions for halos in the three mass bins defined above differ considerably less than the distributions of velocity. Specific angular momentum distribution is thus not a very sensitive probe of halo mass. Indeed, the halo mass is proportional to the second moment of the velocity distribution $\langle v^{2}\rangle$, while Patel et al. (2018) find that the halo mass depends on the median $j$ of satellites as $M_{\rm 200c}\propto j^{1.3}$, and this relation has a large scatter. This motivated us to search for an alternative proxy of $M_{\rm 200c}$ which is both sensitive to the value of the mass and robust against effects of the LMC.

### v2

@@ Abstract
We present a study illustrating the effects of the passage of a Large Magellanic Cloud (LMC) mass satellite on the distance and velocity distributions of satellites in $\Lambda+$Cold Dark Matter simulations of Milky Way (MW) sized halos. In agreement with previous studies, we find that during such a passage the velocity distribution develops a high-velocity tail, which can bias velocity-based virial halo mass estimates. When the velocity distribution of MW satellites is corrected for effects of the LMC passage, it is consistent with the distributions in halos of masses as low as $M_{\rm 200c}=8\times 10^{11}\,\,M_{\odot}$ and as high as $1.5\times 10^{12}\,\,M_{\odot}$. We present a new halo mass estimator $M_{\rm 200c}=c\sigma^{2}_{\rm 3D}r_{\rm med}$, where $c$ is the coefficient calibrated using satellite systems in the simulated MW-sized halos, $\sigma^{2}_{\rm 3D}$ is the variance of 3D velocities taken with the sign of the radial velocity of each satellite, and $r_{\rm med}$ is the median halocentric distance of the satellites. We show that the estimator has only $s=8\%$ scatter around the median relation of the estimated and true halo masses and deviates by $<2s$ from the median during the pericentric passage of an LMC-like subhalo. This is because $\sigma^{2}_{\rm 3D}$ and $r_{\rm med}$ deviate in the opposite directions during such passages. We apply the estimator to the MW satellite system and estimate the virial mass of the Milky Way of $M_{\rm 200c}=9.96\pm 1.45\times 10^{11}\,\,M_{\odot}$, in good agreement with several recent estimates using other methods.
@@ 1 Introduction
The improved velocity accuracy and sophistication of the analyses led to reduced uncertainties in halo mass estimates and reduced discrepancies between different studies. Nevertheless, recent MW halo virial mass measurements still allow a fairly wide range of values $\approx 6-20\times 10^{11}\,M_{\odot}$ (e.g., Deason et al., 2021; Vasiliev et al., 2021; Correa Magnus & Vasiliev, 2022; Koposov et al., 2023; Zhou et al., 2023; Roche et al., 2024).
@@
Such a large LMC mass results in perturbations of velocities and orbits of individual MW satellites (Erkal et al., 2020; Pace et al., 2022; Makarov et al., 2023), significant motions of the inner Milky Way, and asymmetries of the potential in which dwarf satellites and orbit (Gómez et al., 2015; Erkal et al., 2020; Garavito-Camargo et al., 2019; Garavito-Camargo et al., 2021; D’Souza & Bell, 2022). In particular, LMC induces the uniform reflex motion of the Milky Way disk relative to the outer halo and distorts the distribution of velocities. The latter leads to negative mean radial velocity of the tracers, such as halo stars, globular clusters, and satellites.
@@
These effects have been detected using different tracers in the Milky Way halo (Petersen & Peñarrubia, 2020, 2021; Erkal et al., 2021; Conroy et al., 2021; Yaaqib et al., 2024). In particular, LMC both induces the motion of the MW disk relative to the outer halo and affects the velocities of the outer satellites, such as Antlia II (Ji et al., 2021) and Leo I (Erkal et al., 2020). The use of the latter in particular tended to produce higher halo mass estimates in prior analyses. Erkal et al. (2020) also showed that the mass estimator of Watkins et al. (2010) applied to the velocities of satellites and stellar tracers overestimates the enclosed mass $M(<r)$ by up to $50\%$ at large $r$ if the LMC-induced effects are neglected. Likewise, Deason et al. (2021) showed that the motion induced by LMC bias mass estimates based on the 6D phase space information of halo stars. The presence of the bound Magellanic Clouds pair also affects statistical estimates of the Milky Way halo mass (González et al., 2013). Finally, the approach of the LMC has a large effect on the spatial distribution of MW satellites (Nadler et al., 2020; Manwadkar & Kravtsov, 2022; Garavito-Camargo et al., 2023).
@@
Large effects of the LMC on the phase-space distribution of the potential Milky Way tracers need to be accounted for in theoretical inferences of halo mass that use these tracers. This is challenging, however, because only $\approx 5-10\%$ of the MW-sized halos have a pair of LMC-like satellites with a pericenter close to $\approx 50$ kpc (Boylan-Kolchin et al., 2010; Busha et al., 2011). Thus, large samples of halos need to be simulated with a high resolution sufficient to resolve and model satellite evolution reliably. Alternatively, the analyses can focus on the properties of tracers that are not significantly affected by the dynamical perturbations due to the LMC, such as specific angular momentum (Patel et al., 2018).
@@
In this study, we explore the effects of the LMC on the velocity distribution of satellite systems using a Caterpillar suite of high-resolution simulations of MW-sized halos (Griffen et al., 2016, see Section 2). We demonstrate that similarly to the real LMC, LMC-like objects in simulations induce motions onto their host halo and its satellites which creates a high-velocity tail in the distribution of halocentric 3D satellite velocities (Section 3.1). We also show how accounting for this tail affects the conclusion about MW halo mass based on the velocity distribution. We then present a new halo mass estimator that has only $\approx 8\%$ scatter around the true mass when applied to simulated halos (Section 3.2) and which is affected by the LMC-like objects by only $\lesssim 10\%$. We use the estimator calibrated with the simulated halos to infer the MW halo mass of $M_{\rm 200c}=9.93\pm 1.45\times 10^{11}\,M_{\odot}$. We summarize our results and conclusions in Section 4.
@@ 2 Modeling Milky Way satellite system
The mass evolution tracks of subhaloes of MW-sized host haloes are used as input for the GRUMPY galaxy formation model, which evolves various properties of gas and stars of the galaxies they host. It is a regulator-type galaxy formation framework (e.g., Krumholz & Dekel, 2012; Lilly et al., 2013; Feldmann, 2013) designed to model galaxies of $\lesssim L_{\star}$ luminosity (Kravtsov & Manwadkar, 2022) which follows the evolution of several key galaxy properties by solving a system of coupled different equations.
The model accounts for UV heating after reionization and associated gas accretion suppression onto small mass haloes, galactic outflows, a model for gaseous disk and its size, molecular hydrogen mass, star formation, etc. The evolution of the half-mass radius of the stellar distribution is also modeled. The galaxy model parameters used in this study are identical to those used in Manwadkar & Kravtsov (2022). Importantly for this study, the model was shown to reproduce the luminosity function and radial distribution of the Milky Way satellites and the size-luminosity relation of observed dwarf galaxies (Manwadkar & Kravtsov, 2022). In Pham et al. (2023) we also showed that the flattening of the subhalo distribution in the Caterpillar suite is consistent with observed flattened distribution of Milky Way satellites. Here we use the model to predict luminosities and stellar masses of satellite galaxies around the MW-sized haloes from the Caterpillar suite.
@@
Figure 3: The cumulative distribution of 3D velocities of satellites in the Milky Way and simulated halos of different halo mass ranges. The velocity distribution of the Milky Way satellites is shown by the thick solid line. The long-dashed, dot-dashed, and dashed lines show velocity CDFs of satellites in the low, intermediate, and high mass host halo mass ranges. The satellites have been selected using the distrance distribution matching procedure described in the text (Section 3.1).
Figure 4: Top panel: evolution of distance (in physical kpc) of two LMC-like satellites to the center of their host halo from $z=0.5$ to $z=0$ in two different Caterpillar host halos. The satellites are on their first infall and reach their pericenter of $\approx 40-50$ kpc at $z_{\rm LMC,peri}\approx 0.04-0.05$. The dotted lines show evolution of the virial radii $R_{\rm 200c}(z)$ of their host halos. The upper $x$-axis shows the lookback time corresponding to the redshifts shown in the bottom $x$-axis. Bottom panel: Evolution of halo masses (solid lines) and stellar masses (dashed lines) of the two LMC-like satellites shown in the top panel. The satellites reach peak halo masses of $M_{\rm peak}\approx 1.16\times 10^{11}\,M_{\odot}$ and $M_{\rm peak}\approx 1.35\times 10^{11}\,M_{\odot}$ at $z\approx 0.1$ before they start to lose mass due to tidal stripping. The stellar mass is computed using our galaxy formation model and continues to grow due to ongoing star formation and reaches $M_{\star}\approx 3\times 10^{9}\,M_{\odot}$, comparable to the stellar mass of the observed LMC.
@@
3 Results
3.1 Effect of the LMC on the velocity distribution of satellites
Comparison to the Milky Way requires a satellite population whose cumulative Galactocentric distance distribution matches that of the MW satellites. To obtain this, we draw a set of random distances from the MW satellite distance distribution using the inverse transform sampling method and spline approximation of the inverse cumulative distribution function of MW satellite distances. We then selected subhalos in each simulated MW-sized halo that were closest to each of the drawn distances. Figure 1 shows that this distance-matching technique produces a satellite population with cumulative distance distribution consistent with the distance distribution of MW the satellites. Note that both here and below the distance $r$ is the distance to the center of the host halo defined by the halo finder (corresponding to the central density peak). At the same time, Figure 2 demonstrates that the 3D velocity distribution of the distance-matched subhalos is statistically consistent with the velocity distribution of Milky Way satellites.
We also checked the effect of distance-matching on the cumulative distributions of satellite velocity’s radial and tangential components. Distance matching generally slightly worsens the agreement of the model and observed radial velocity distributions, but the effect of the matching on the radial velocity distribution is small. Distance matching does improve agreement of the tangential velocity distributions and it is this improvement that is responsible for the better match of the 3d velocity distribution.
Hosts 4 and 53 from the Caterpillar host halo tracks include the passage of subhalos that reach peak masses of $M_{\rm peak}=1.16\times 10^{11}\,\,M_{\odot}$ and $M_{\rm peak}=1.35\times 10^{11}\,\,M_{\odot}$, respectively. These are comparable to the mass of the LMC with $M_{\rm peak}=1.3\pm 0.3\times 10^{11}\,M_{\odot}$ (Erkal et al., 2019; Vasiliev et al., 2021; Shipp et al., 2021; Watkins et al., 2024). The time evolution of the halocentric distance and mass of these subhalos are shown in Figure 4. The satellites reach their peak masses $M_{\rm peak}\approx 1.16\times 10^{11}\,M_{\odot}$ and $M_{\rm peak}\approx 1.35\times 10^{11}\,M_{\odot}$ at $z\approx 0.1$ and begin to lose mass due to tidal stripping during their pericentric passage at $z\approx 0.05$ or about $\approx 0.7$ Gyr after they cross $R_{\rm 200c}$ of their host halo. This is similar to the estimates of the accretion time of the LMC in the Milky Way (Petersen & Peñarrubia, 2020). As the LMC-like satellites continue to be tidally stripped, stellar mass steadily increases due to continued star formation and reaches a peak of $M_{\star}\approx 3\times 10^{9}\,M_{\odot}$, comparable to the observational estimates of the LMC’s stellar mass.
@@ 3.1 Effect of the LMC on the velocity distribution of satellites
For one of the host halos, Figure 5 shows the satellite velocity distribution at $z=0$ and at the time of pericentric passage $z\approx 0.05$, as well as the 15.9 - 84.1 and 2.3 - 97.7 percentile intervals of the velocity distribution in bins of distance. The effect of the LMC-like satellite is greatest near the pericenter, where the subhalo velocity distribution becomes skewed toward larger negative velocities. At $z=0$, the velocities return to a symmetric distribution. The effect of the LMC-like subhalo passage on the cumulative velocity distribution is shown in Figure 6, which shows that the satellite distribution has a tail of higher velocities near the pericentric passage. The cumulative velocity distribution at $z=0$, on the other hand, is similar to the distribution before the LMC-like subhalo entered the host halo at $z=0.1$.
The lower panel of Figure 5 shows that the cumulative distance distribution of satellites becomes more centrally concentrated during the pericentric passage of the LMC-like subhalo. Although the physical mechanism of the distance distribution change warrants a detailed investigation, it is most likely a result of contraction of the satellite orbits in response to the deepening of the potential due to the LMC mass when it is near the pericenter. Although the LMC pericenter is $\approx 50$ kpc, the virial radius of the LMC and its mass is about 100 kpc and encompasses and affects a significant fraction of the Milky Way’s virial volume. Like the velocity distribution, the distance distribution at $z=0$ becomes closer to the distribution before the passage at $z=0.1$, although the distribution at $r\gtrsim 180$ kpc remains somewhat different most likely due to the presence of the LMC counterpart at $r=150$ kpc at $z=0$ (see Figure 4).
@@
The physical origin of the large change in the velocity and distance distribution of satellites during the pericentric passage of an LMC-like subhalo is the reflex motion of the central region of the MW-sized host halo and associated compression of satellite distribution and velocities in response to the gravitational pull of the massive subhalo and due to a deepening of the potential when the LMC counterpart reaches the pericenter of its orbit.
We have computed both the overall mean 3D velocity of reflex motion of the simulated satellites relative to the halo center ($\mathbf{v}_{\rm travel}$ in Petersen & Peñarrubia, 2020, 2021) and the mean radial velocity of satellites after the $x$, $y$, and $z$ components of the mean 3D velocity are subtracted from the velocity components of the satellites. The reflex motion velocity ($v_{\rm travel}$) for the two host halos with LMC-like satellites is 44 and 38 km/s. The mean radial velocities after the mean velocity subtraction are $\langle v_{r}\rangle=-24$ km/s and $-44$ km/s. These numbers are comparable to the corresponding mean 3D velocity of motion of the inner Milky Way and mean radial velocity of satellites estimated by Petersen & Peñarrubia (2021, see their Table 1). Note that the mean negative radial velocity of satellites reflects a distortion of their velocity distribution by the LMC in addition to the overall mean reflex motion. Such distortion (dubbed “compression” by the authors) is now reliably detected in the velocity tracers of the Milky Way (Petersen & Peñarrubia, 2021; Yaaqib et al., 2024). We find that in simulations by $z=0$ when LMC counterparts move to $r\approx 150$ kpc, both the reflex motion velocity and the mean radial velocity decrease by a factor of three.
@@
The large effect of the LMC on the velocity and distance distribution needs to be taken into account in the estimates of MW halo mass (Erkal et al., 2020). Although the effect becomes small well after the passage, the LMC in the Milky Way has just passed its pericenter and therefore effects on the MW satellite velocities are expected to be significant. To illustrate the effect of change in velocity distribution on halo mass inference, we approximate the difference in the velocity distributions at the pericentric passage of the LMC-like subhalo and at $z=0$ as a function of 3D velocity of simulated satellites using a cubic spline. We then use the spline to compute the correction for each satellite of the Milky Way using its 3D velocity. Note that this correction is estimated using only two MW-sized host halos with LMC-like satellites that are present in the Caterpillar sample and is thus uncertain. We use it here for illustration of the potential effects of the LMC on the virial mass inference.
@@
Patel et al. (2018, see also ) used the specific angular momenta of satellites to estimate MW halo mass and showed that they are not sensitive to deviations from dynamic equilibrium. The upper panel of Figure 9 shows the cumulative distribution of specific angular momentum $j=|\mathbf{r}\times\mathbf{v}_{\rm 3D}|$ at three different epochs, $z=0.1$, $z=z_{\rm LMC,peri}$ and $z=0$, for one of the host halos with an LMC-like satellite passing through its pericenter at $z_{\rm LMC,peri}\approx 0.05$. The effect of the LMC-like satellite on the $j$ distribution is indeed very small, which is consistent with the conclusion of Patel et al. (2018). Thus, the use of the angular momentum to estimate halo mass would not require a correction for the effect of the LMC.
At the same time, the lower panel of Figure 9 shows that the cumulative $j$ distributions for halos in the three mass bins are not sufficiently different to derive an accurate mass estimate. Patel et al. (2018) find that the median specific angular momentum, $j_{\rm med}$ of satellites in halos scales as $j_{\rm med}\propto M_{\rm 200c}^{0.78\pm 0.03}$. Given the halo mass range of the Caterpillar host halo sample, we would expect $j_{\rm med}$ to differ at least by a factor of $\approx 1.5$ between the LM and HM mass bins. However, the median in the lower panel of Figure 9 varies less than this and there is little difference in the distributions of $j$ for the IM and HM samples. We think this is most likely due to the wide span of $j$ distribution for hosts of a given mass. When dealing with a finite number of host halos and their satellite, specific subsamples from this wide distribution can result in a large scatter of $j_{\rm med}$ that obscures the average trend of $j_{\rm med}$ with $M_{\rm 200c}$.
These considerations indicate that both velocity and specific angular momentum distributions quantified using small samples of simulated halos do not allow an accurate estimate of MW halo mass. This motivated us to search for an alternative proxy of $M_{\rm 200c}$ which is both sensitive to the value of the mass and robust against effects of the LMC.